OpenStudy (anonymous):

Solve the system using elimination. x = –4 + 3y 12y = 21 + 3x Can someone please go step by step for me? Thank you

OpenStudy (anonymous):

y = 3 and x = 5

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

in elimination method first make the coefficients of any two variables same by mutiplying a number, then add or subtract themm to eliminate a variable. last get the values.

OpenStudy (anonymous):

to solve simultaneously you put them in the form x + y = number that is, put the variables like x and y on the left hand side and the constant on the right. so in this case it would look like:--> x -3y = -4 3x -12y = -21 after this you need to get rid or/eliminate one of the variables...i'll eliminate x so i'll be left with y. since the x terms need to be the same to be elimated, multiply the top eq. by 3 so that you get: 3x -9y = -12 - 3x -12y = -21 _____________________ then atter subtracting those two eqs you get 3y = 9 and y = 3 then substitute back into the original eq: x - 3(3) = -4 so : x -9 = -4 and x = 5

OpenStudy (anonymous):